OpenStudy (anonymous):

jeremiah has a long term savings plan. For 10 years , he has been investing $150 a month, earning 4.25 interest compounded monthly. How much more would he have saved if he had chosen to make deposits of @200 a month?

OpenStudy (anonymous):

\[-\frac{D \left(1-(1+R)^N\right)}{R}\]

OpenStudy (anonymous):

\[(200-150)\left(-\frac{ \left(1-(1+.0425)^{120}\right)}{.0425}\right)\] \[172487.\]